# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v1

@@ 2   Related Work
A landmark 2016 paper introduced a new paradigm to design iterative methods for computing $\operatorname*{polar}(\bm{M})$ [34]. We describe this paradigm in more detail in Section 4, but the main insight is as follows. Padé methods choose the update rule to be an approximation to $\operatorname*{sign}(x)$ of a given degree that is optimally accurate in the neighborhood of $x=1$. Instead, we should choose the approximation to $\operatorname*{sign}(x)$ that is optimal over an interval $[\ell,1]\subset\mathbb{R}_{\geq 0}$ that contains the singular values. Moreover, after each step of the algorithm, the range of the singular values changes; therefore, we adapt the update rule at each iteration to match the new lower bound $\ell$. When the range of the singular values is large, this approach ensures that the update rule shrinks it as quickly as possible. As the algorithm proceeds and the interval shrinks to a small neighborhood of $1$, the update rule approaches that of a Padé method, maintaining the same high order of convergence as it has. Within the class of odd rational functions whose numerators and denominators have degree $12q+$ 1 and $2q$, respectively, an explicit formula for this optimal approximation to $\operatorname*{sign}(x)$ on any interval $[\ell,1]$ was found by Zolotarev. For $q=1$, this function coincides exactly with the updates of the dynamically weighted Halley’s method referred to above. It is shown that these rationals have remarkable convergence properties for any $q$. For even faster convergence than QDWH, the zolo-pd method chooses $q=17$ [34]. Finally, these methods admit the same QR-based implementation trick as QDWH.
@@
The designers of Muon realized that, due to the extreme efficiency requirements and lax accuracy requirements of their setting, rational-based methods from the numerical analysis literature are inapplicable. However, polynomial-based iteration schemes can take full advantage of GPUs because they use only matrix-matrix products in half-precision arithmetic, not inverses or QR decompositions. The preference for speed over accuracy motivates methods that aim to quickly produce coarse approximations, even at the cost of asymptotic convergence. Examples include the proposals of Jordan [23] and You [43, 10], as discussed in Section 1.2. Like Chen and Chow [12], Jordan found that convergence in the initial phase can be accelerated by choosing update rules that have a large derivative near zero, so as to increase the small singular values as much as possible at each iteration. You furthermore chose to use different update rules at each iteration, allowing extra flexibility to tune the trade-off between speed and accuracy. Both used degree-5 polynomials that were found through gradient descent on heuristic objective functions. These proposals were compared to Newton-Schultz [^2] Jordan[23] actually compares to $2x-\frac{3}{2}x^{3}+\frac{1}{2}x^{5}$, whereas the true degree-5 Newton-Schulz iteration is $(15x-10x^{3}+3x^{5})/8$. However, the difference in performance is negligible for the first few iterations., but not to Chen and Chow’s method. We find that our method outperforms them all.
@@ 3   Approximations by compositions of polynomials
To derive a GPU-friendly method for computing $\operatorname*{polar}(\bm{M})$, we limit ourselves to the following GPU-friendly operations:
@@
Hence, for an odd polynomial $p(x)=a_{0}x+a_{1}x^{3}+\cdots a_{q}x^{2q+1}$ we can compute
@@
Letting $\ell$ and $u$ be the minimum and maximum singular values of $\bm{M}$, respectively, our goal is now to find an odd polynomial $p$ so that $p(x)$ is as close to $1$ as possible on $[\ell,u]$. Indeed, given that $p(\bm{M})-\operatorname*{polar}(\bm{M})=\bm{U}(p(\bm{\Sigma})-\bm{I})\bm{V}^{\mathsf{T}}$, by the unitary invariance of the spectral norm we have that:
$\|p(\bm{M})-\operatorname*{polar}(\bm{M})\|_{2}\leq\max_{x\in[\ell,u]}|p(x)-1|.$ (4)
@@
By the unitary invariance of the spectral norm, we have that (7) is equivalent to
@@
For completeness, the equivalence is verified in Appendix C. In other words, the problem given in (7) reduces to that of finding a “uniform” or “minimax” approximation to the constant function $x\mapsto 1$ over the interval $[\ell,u]$, as given in (8). Uniform approximation on an interval by polynomials or rational functions of a given degree is a central topic in approximation theory; see e.g. [45]. Here, we seek an approximation of a particular form—a composition of odd polynomials of fixed degrees. In the next section, we solve the optimizing problem of (8) and use the solution to create Polar Express.
@@ 4.1   Greedy is optimal
Our key observation is that the polynomial used in each step can be chosen greedily, given the choice of polynomials from the previous iterations. For the first iteration, we choose $p_{1}$ so as to map the interval $[\ell,u]$ as close to $1$ as possible. That is, it minimizes $\max_{x\in[\ell,u]}|1-p_{1}(x)|$. The image of $p_{1}$ will be a new interval $[\ell_{2},u_{2}]$, where
@@
We now pick $p_{2}$ to map the interval $[\ell_{2},u_{2}]$ as close to $1$ as possible, obtaining a new interval $[\ell_{3},u_{3}]$ that is the image of $[\ell,u]$ through $p_{2}\circ p_{1}$. We continue this process for as many iterations as desired. This same idea was used in zolo-pd [34] to approximate $\operatorname*{polar}(\bm{M})$ by a composition of rational functions; here we use a composition of polynomials.
(a) The left figure compares the resulting composition (for $T=6$ and $d=5$) of polynomials given by Polar Express ($\ell=0.001$), You’s method (which is defined up to 6 iterations), Newton-Schulz, and Jordan’s method for approximating $\operatorname*{sign}(x)$. The right figure demonstrates error convergence of the methods. Note the slow initial convergence of the Newton-Schulz method.
Figure 2:
The following theorem guarantees that this process finds the solution to (8), and thereby also (7). The scheme is also outlined in Figure 2, which demonstrates the evolution of the lower bounds $\ell_{t}$, the upper bounds $u_{t}$, and the polynomials $p_{t}$ throughout the iterations.
@@
Our theory translates entirely to this more general case. However, for simplicity we assume $d=d_{t}$ for all $t=1,\ldots,T$. Our setting is similar to that of [29], which considers the closely related problem of choosing the depth $T$ and degrees $d_{1},\ldots,d_{T}$ such that $p$ approximates $\operatorname*{sign}$ up to a prescribed error tolerance while minimizing the number of scalar multiplications. Interestingly, from [29, Table 2] the optimal choice of degrees is $d_{t}=5$ for almost all iterations. This justifies choosing $d$ to be a constant and our use of $d=5$ in particular.
@@
From optimality guarantee of Theorem 4.1, we know that our method converges at least as fast as the Newton-Schulz iteration of the same degree. Combining this fact with an existing analysis of Newton-Schulz, we immediately get the following convergence guarantee showing that our method enjoys faster than exponential convergence.
@@
In fact, Theorem 4.3 underestimates how fast our method converges. For degree $d=5$, our method converges about twice as fast as Newton-Schulz (cf. [12, Section 3.1]). Furthermore, the same analysis applies even if $p^{*}$ is constructed using alower bound $\ell$ that is , where $\ell^{\prime}\geq\sigma_{\min}(\bm{M})\geq\ell$. Intuitively, when $\ell=u$, the polynomial $p^{*}$ is coincides exactly with the Newton-Schulz method. Mistakenly setting $\ell$ in the range $(\sigma_{\min},u)$, we obtain a polynomial whose convergence speed is somewhere between that of the optimal polynomial and Newton-Schulz, so the guarantee of Theorem 4.3 can be extended to it as well (cf. [12, Theorem 3.3]).
@@ 4.2   Finding the optimal polynomial for each iteration
Theorem 4.1 shows that we can solve (8) by greedily choosing the optimal approximation $p_{t}\in\mathbb{P}_{d}^{\operatorname*{odd}}$ for each interval $[\ell_{t},u_{t}]$ for $t=1,\ldots,T$. In this section, we show how to to find each $p_{t}$. Since we are now focused on just one iteration, we drop the subscripts. Given $\ell$ and $u$, we wish to solve the following optimization problem:
@@
That is, we seek a minimax or uniform approximation of the function $x\mapsto 1$ on $[\ell,u]$ from the set of odd polynomials. (Equivalently, we seek a minimax optimal approximation to $\operatorname*{sign}(x)$ on $[-u,-\ell]\cup[\ell,u]$.) Problems of this form are well-studied in approximation theory and numerical analysis. The key mathematical insight underlying the solution is the Equioscillation Theorem, which we state formally for our setting in Lemma A.1. This theorem gives a surprising characterization of the optimal approximant: an odd $p$ is optimal for degree $12q+$ 1 if and only if there is a set of $q+2$ equioscillating points. These are points at which $p$ achieves its maximum approximation error ±$\pm E$, and for which the sign of the error alternates. Even if $E$ is not known in advance, finding a set of $q+2$ equioscillating points for a given $E$ serves as a certificate that no better approximation error is achievable. The Equioscillation Theorem is the basis of the Remez algorithm [38, 39], a general tool that can be used to find (nearly) optimal polynomial approximations of a given of any degree to any function on any interval. With very minor modifications to handle the constraint that $p$ be odd, Remez can be used to directly solve (15). However, the Remez algorithm is opaque, complex, and difficult to implement correctly. Fortunately, we do not need the Remez algorithm in its full generality to solve our problem. We seek only low degree approximants, and the function we wish to approximate is a constant. For $d=3$, we can actually derive an explicit, closed form solution to (15) using the Equioscillation Theorem. Up to rescaling, the optimal polynomial turns out to be the same one derived in Chen and Chow by different means [12]. For degree $d=5$, we present Algorithm 2, a much simpler way of solving (15) that is mathematically equivalent to Remez in our setting. This algorithm is implemented in its entirety in Appendix G.
@@
We briefly describe the solution for $d=3$. We seek a polynomial of the form $p(x)=ax+bx^{3}$. The Equioscillation Theorem stipulates that $p$ must have an equioscillating set of size 3. For $p$ to achieve its maximum error at a point $x$, $x$ must be a local extremum of $p(x)-1$ on the interval $[\ell,u]$. Thus, for $x$ to be eligible for membership in the equioscillating set, it must either be a true local extremum of $p(x)-1$ that lies in $[\ell,u]$, or else one of the endpoints $\ell,u$. However, because $p$ is an odd cubic, it has at most one true local extremum on $\mathbb{R}_{\geq 0}$. Thus, for build an equioscillating set of three points, we must include $\ell,u$ and $p$’s unique positive local extreumum. This local extremum of $p$ occurs at $\sqrt{\frac{-a}{3b}}$. Therefore, we seek $a,b$ such that
$p(\ell)=1-E\qquad p\left(\sqrt{\frac{-a}{3b}}\right)=1+E\qquad p(u)=1-E$ (16)
@@
for some unknown $E$. This is a system of three equations in three variables. The solution $p(x)=ax+bx^{3}$ is most easily expressed as follows. Let $p_{\operatorname*{NS}}(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$. Then
@@
$p(x)=\beta p_{\operatorname*{NS}}(\alpha x),\text{ where }\alpha=\sqrt{\frac{3}{u^{2}+lu+\ell^{2}}}\text{ and }\beta=\frac{4}{2+\ell u(\ell+u)\alpha^{3}}.$
@@
We now turn to the degree-5 case. The intuition of Algorithm 2 is as follows. For any fixed set of four points $\ell<q<r<u$, we can find an degree-5 odd polynomial $p$ that satisfies
$p(\ell)=1-E\quad p(q)=1+E\quad p(r)=1-E\quad p(u)=1+E$
@@
for some $E$ by solving a $4\times 4$ linear system. Likewise, for any fixed degree-5 odd $p$, we can find its four (or fewer) local extrema on $[\ell,u]$ as follows: they occur at $\ell,u$ and the roots of $p^{\prime}$, which is an even degree-4 polynomials whose roots can be found by the quadratic formula. Algorithm 2 simply alternates between these two steps until the points $q,r$ converge. Once they have converged, they form an equioscillating set, so $p$ is the optimal polynomial. For more details, please see Appendix D.
@@ 4.3   Upper and lower bounds on the singular values
To instantiate our method, we need upper and lower bounds $u$ and $\ell$ on the singular values of the input matrix $\bm{M}$. A trivial upper bound is given by $\|\bm{M}\|_{\text{F}}$. For $\bm{M}\in\mathbb{R}^{m\times n}$ with $n\leq m$, this can overestimate $\sigma_{\max}(\bm{M})$ by a factor of $\sqrt{n}$ in the worst case. However in practice, the gradient matrices of the weights of dense linear layers in neural networks tend to have small effective rank [46]. Consequently, the Frobenius norm tends to be a reasonably good bound on the spectral norm that is loose only by a small constant factor. For stability and consistency, we rescale the input matrix by setting $\bm{X}_{0}=\bm{M}/\|\bm{M}\|_{\text{F}}$ and $u=1$. This rescaling has no effect on the error of our method.
It is difficult to efficiently find a good lower bound on the smallest singular value, so we are forced to guess. Fortunately, the consequences of a bad guess are not severe. As discussed above, the method will eventually converge for any $\ell\in(0,u]$, and even an order of magnitude error only delays convergence by a few iterations. For matrices stored in floating point arithmetic, the singular values are usually larger than machine precision $\epsilon_{\text{mach}}$ [8], so a good guess is to set $\ell\approx\epsilon_{\text{mach}}$. In our numerical experiments we work in bfloat16, hence we set $\ell=10^{-3}$ and $u=1$. Since we use these bounds for all input matrices, we can precompute the optimal polynomials once and apply them to as many inputs as we want.

### v2

@@ 2   Related Work
A landmark 2016 paper introduced a new paradigm to design iterative methods for computing $\operatorname*{polar}(\bm{M})$ [33]. We describe this paradigm in more detail in Section 4, but the main insight is as follows. Padé methods choose the update rule to be an approximation to $\operatorname*{sign}(x)$ of a given degree that is optimally accurate in the neighborhood of $x=1$. Instead, we should choose the approximation to $\operatorname*{sign}(x)$ that is optimal over an interval $[\ell,1]\subset\mathbb{R}_{\geq 0}$ that contains the singular values. Moreover, after each step of the algorithm, the range of the singular values changes; therefore, we adapt the update rule at each iteration to match the new interval. When the range of the singular values is large, this approach ensures that the update rule shrinks it as quickly as possible. As the algorithm proceeds and the interval shrinks to a small neighborhood of $1$, the update rule approaches that of a Padé method, maintaining the same high order of convergence as it has.
Within the class of odd rational functions whose numerators and denominators have degree $12q+$ 1 and $2q$, respectively, an explicit formula for this optimal approximation to $\operatorname*{sign}(x)$ on any interval $[\ell,1]$ was found by Zolotarev. It was shown that these rationals have remarkable convergence properties for any $q$ [33]. For $q=1$, this optimal approximation coincides exactly with the dynamically weighted Halley’s method (QDWH) referenced above. For even faster convergence than QDWH, [33] proposed the zolo-pd method, which uses $q=17$. Finally, these methods all admit the same QR-based implementation trick as QDWH.
Adaptive polynomial methods.
@@
Application within Muon.
The designers of Muon realized that, due to the extreme efficiency requirements and lax accuracy requirements of their setting, rational-based methods from the numerical analysis literature are inapplicable. However, polynomial-based iteration schemes can take full advantage of GPUs because they use only matrix-matrix products in half-precision arithmetic, not inverses or QR decompositions. The preference for speed over accuracy motivates methods that aim to quickly produce coarse approximations, even at the cost of asymptotic convergence. Examples include the proposals of Jordan [22] and You [44, 9], as discussed in Section 1.2. Like Chen and Chow [11], Jordan found that convergence in the initial phase can be accelerated by choosing update rules that have a large derivative near zero, so as to increase the small singular values as much as possible at each iteration. You furthermore chose to use different update rules at each iteration, allowing extra flexibility to tune the trade-off between speed and accuracy. Both used degree-5 polynomials that were found through gradient descent on heuristic objective functions. These proposals were previously compared to Newton-Schultz [^3] Jordan [22] actually compares to $2x-\frac{3}{2}x^{3}+\frac{1}{2}x^{5}$, whereas the true degree-5 Newton-Schulz polynomial is $(15x-10x^{3}+3x^{5})/8$. However, the difference in performance is negligible for the first few iterations., but never to Chen and Chow’s method. We find that our method outperforms them all.
@@ 3   Approximations by Compositions of Polynomials
To design a GPU-friendly method for computing $\operatorname*{polar}(\bm{M})$, we limit ourselves to the following GPU-friendly operations:
@@
Hence, for an odd polynomial $p(x)=a_{0}x+a_{1}x^{3}+\cdots+a_{q}x^{2q+1}$ we can compute
@@
Given that $\operatorname*{polar}(\bm{M})-p(\bm{M})=\bm{U}(\bm{I}-p(\bm{\Sigma}))\bm{V}^{\mathsf{T}}$, and by the unitary invariance of the spectral norm, we have that (6) is equivalent to
@@
For completeness, the equivalence between (6) and (7) is proven in Appendix C.
(a) The left figure compares the composition (for $T=6$ and $d=5$) of polynomials given by Polar Express ($\ell=0.001$), You’s method (which is defined up to 6 iterations), Newton-Schulz, and Jordan’s method for approximating $\operatorname*{sign}(x)$. The right figure demonstrates the convergence of the methods on $[0.001,1]$. Note the slow initial convergence of Newton-Schulz.
Figure 2:
In other words, the problem given in (6) reduces to that of finding a “uniform” or “minimax” approximation to the constant function $x\mapsto 1$ over the interval $[\ell,u]$, as given in (7). Uniform approximation on an interval by polynomials or rational functions of a given degree is a central topic in approximation theory; see e.g. [46]. Here, we seek an approximation of a particular form—a composition of odd polynomials of fixed degrees. In the next section, we solve the optimization problem of (7) and use the solution to create Polar Express. Figure 2 (a) shows the resulting $p^{*}$ polynomial labeled as PolarExp, as compared to the Jordan’s method in [22], and the six iterations of You’s method in [9].
@@ 4.1   Greedy is optimal
The key observation is that the polynomial used in each iteration can be chosen greedily, given the choice of polynomials from the previous iterations. For the first iteration, we choose $p_{1}$ so as to map the interval $[\ell,u]$ as close to $1$ as possible. That is, it minimizes $\max_{x\in[\ell,u]}|1-p_{1}(x)|$. The image of $p_{1}$ will be a new interval $[\ell_{2},u_{2}]$, where
@@
We now pick $p_{2}$ to map the interval $[\ell_{2},u_{2}]$ as close to $1$ as possible, obtaining a new interval $[\ell_{3},u_{3}]$ that is the image of $[\ell,u]$ through $p_{2}\circ p_{1}$. We continue this process for as many iterations as desired.
The following theorem guarantees that this process finds the solution to (7), and thereby also (6). The scheme is also outlined in Figure 2 (b), which demonstrates the evolution of the lower bounds $\ell_{t}$, the upper bounds $u_{t}$, and the polynomials $p_{t}$ across iterations.
@@
Our theory translates entirely to this more general case. However, for simplicity we assume $d=d_{t}$ for all $t=1,\ldots,T$. Our setting is similar to that of [28], which considers the closely related problem of choosing the depth $T$ and degrees $d_{1},\ldots,d_{T}$ such that $p$ approximates $\operatorname*{sign}$ up to a prescribed error tolerance while minimizing the number of scalar multiplications. Interestingly, from [28, Table 2] the optimal choice of degrees is $d_{t}=5$ for almost all iterations. This justifies choosing $d$ to be a constant and our use of $d=5$ in particular.
@@
From the optimality guarantee of Theorem 4.1, we know that our method converges at least as fast as the Newton-Schulz iteration of the same degree. Combining this fact with an existing analysis of Newton-Schulz, we immediately get the following convergence guarantee showing that our method enjoys faster than exponential convergence.
@@
In fact, Theorem 4.3 underestimates how fast our method converges. For degree $d=5$, our method converges about twice as fast as Newton-Schulz (compare with [11, Section 3.1]). Furthermore, the same analysis applies even if $p^{*}$ is constructed using a “lower bound” $\ell$ that was too high. That is, replacing $\ell$ on the right-hand side of (13) by $\sigma_{\min}$, the theorem holds even if $p^{*}$ is constructed to be optimal on the interval $[\ell,1]$ for $\ell>\sigma_{\min}$. Intuitively, when $\ell=u=1$, the polynomial $p^{*}$ coincides exactly with the Newton-Schulz method. Mistakenly setting $\ell>\sigma_{\min}$, we obtain a polynomial that converges slower than the optimal polynomial but faster than Newton-Schulz, so the guarantee of Theorem 4.3 still holds (cf. [11, Theorem 3.3]).
@@ 4.2   Finding the optimal polynomial for each iteration
Theorem 4.1 shows that we can solve (7) by greedily choosing the optimal approximation $p_{t}\in\mathbb{P}_{d}^{\operatorname*{odd}}$ for each interval $[\ell_{t},u_{t}]$ for $t=1,\ldots,T$. In this section, we show how to find each $p_{t}$. Since we are now focused on just one iteration, we drop the subscripts. Given $\ell$ and $u$, we wish to solve the following optimization problem:
@@
That is, we seek a minimax or uniform approximation of the function $x\mapsto 1$ on $[\ell,u]$ from the set of odd polynomials. (Equivalently, we seek a minimax optimal approximation to $\operatorname*{sign}(x)$ on $[-u,-\ell]\cup[\ell,u]$.)
@@
Problems of the form (14) are well-studied in approximation theory and numerical analysis. The key mathematical insight underlying the solution is the Equioscillation Theorem, which we state formally for our setting in Lemma A.1. This theorem gives a surprising characterization of the optimal solution of (14): an odd $p$ is optimal for degree $12q+$ 1 if and only if there is a set of $q+2$ equioscillating points. This is a set of points at which $p$ achieves its maximum approximation error ±$\pm E$, and for which the sign of the error alternates. Even if the optimal approximation error $E$ is not known in advance, finding a set of $q+2$ equioscillating points for a given $E$ serves as a certificate that no better approximation error is achievable. The Equioscillation Theorem is the basis of the Remez algorithm [37, 38], a general tool that can be used to find (nearly) optimal polynomial approximations of a given degree to any function on any interval. With very minor modifications to handle the constraint that $p$ be odd, Remez can be used to directly solve (14).
However, the Remez algorithm is opaque, complex, and difficult to implement correctly. Fortunately, we do not need the Remez algorithm in its full generality to solve our problem. We seek only low degree polynomials, and the function we wish to approximate is the constant function $f(x)\equiv 1$. For $d=3$, we can derive an explicit, closed form solution to (14) using the Equioscillation Theorem. Up to rescaling, the optimal polynomial turns out to be the same one derived in Chen and Chow by different means [11]. For degree $d=5$, we present Algorithm 3, a much simpler way of solving (14) that is mathematically equivalent to Remez in our setting. This algorithm is implemented in its entirety in Appendix G.
@@
We briefly describe the solution for $d=3$. We seek a polynomial of the form $p(x)=ax+bx^{3}$. The Equioscillation Theorem stipulates that $p$ must have an equioscillating set of size 3. For $p$ to achieve its maximum error at a point $x$, $x$ must be a local extremum of $p(x)-1$ on the interval $[\ell,u]$. Thus, for $x$ to be eligible for membership in the equioscillating set, it must either be a true local extremum of $p(x)-1$ that happens to lie in $[\ell,u]$, or else one of the endpoints $\ell,u$. However, because $p$ is an odd cubic, it has at most one true local extremum on $\mathbb{R}_{\geq 0}$. Thus, to build an equioscillating set of three points, we must include $p$’s unique positive local extremum and both endpoints. This local extremum of $p$ occurs at $\sqrt{\frac{-a}{3b}}$. Therefore, we seek $a,b$ such that
$p(\ell)=1-E,\qquad\qquad p\left(\sqrt{\frac{-a}{3b}}\right)=1+E,\qquad\qquad p(u)=1-E$ (15)
@@
for some $E$. This is a system of three equations in three variables. The solution $p(x)=ax+bx^{3}$ is most easily expressed as follows. Let $p_{\operatorname*{NS}}(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$. Then
@@
$p(x)=\beta p_{\operatorname*{NS}}(\alpha x),\quad\text{ where }\alpha=\sqrt{\frac{3}{u^{2}+lu+\ell^{2}}}\quad\text{ and }\quad\beta=\frac{4}{2+\ell u(\ell+u)\alpha^{3}}.$
@@
We now turn to the degree-5 case. The intuition of Algorithm 3 is as follows. For any fixed set of four points $\ell<q<r<u$, we can find an degree-5 odd polynomial $p(x)=ax+bx^{3}+cx^{5}$ that satisfies
$p(\ell)=1-E,\qquad p(q)=1+E,\qquad p(r)=1-E,\qquad p(u)=1+E$
@@
for some $E$ by solving a $4\times 4$ linear system in $a,b,c$ and $E$. Likewise, for any fixed degree-5 odd $p$, we can find its four (or fewer) local extrema on $[\ell,u]$ as follows: they occur at $\ell,u$ and the roots of $p^{\prime}$, which is an even degree-4 polynomial whose roots can easily be found by the quadratic formula. Algorithm 3 simply alternates between these two steps (solving for $a,b,c,E$ and solving for $q,r$) until the points $q,r$ converge. Once they have converged, $\{\ell,q,r,u\}$ forms an equioscillating set, so $p$ is the optimal polynomial. For more details, please see Appendix D.
@@ 4.3   Upper and lower bounds on the singular values
To instantiate our method, we need upper and lower bounds $u$ and $\ell$ on the singular values of the input matrix $\bm{M}$. A trivial upper bound is given by $\|\bm{M}\|_{\text{F}}$. For $\bm{M}\in\mathbb{R}^{m\times n}$ with $n\leq m$, this can overestimate $\sigma_{\max}(\bm{M})$ by a factor of $\sqrt{n}$ in the worst case. However in practice, the gradient matrices of the weights of dense linear layers in neural networks tend to have small effective rank [48]. Consequently, the Frobenius norm tends to be a reasonably good bound on the spectral norm that is loose only by a small constant factor. We rescale the input matrix by setting $\bm{X}_{0}=\bm{M}/\|\bm{M}\|_{\text{F}}$ so that $u=1$.
It is difficult to efficiently find a good lower bound on the smallest singular value, so we are forced to guess. Fortunately, the consequences of a bad guess are not severe. As discussed above, the method will eventually converge for any $\ell\in(0,u]$, and even an order of magnitude error in our guess of $\ell$ only delays convergence by a few iterations. For matrices stored in floating point arithmetic, the singular values are usually larger than machine precision $\epsilon_{\text{mach}}$ [7], so a good guess is to set $\ell\approx\epsilon_{\text{mach}}$. In our numerical experiments we work in bfloat16 where $\epsilon_{\text{mach}}=2^{-7}=0.0078125.$ Hence we set $\ell=10^{-3}$ and $u=1$. Since we use these bounds for all input matrices, we can precompute the optimal polynomials once and apply them to as many inputs as we want.